### v2

# AdaGC: Enhancing LLM Pretraining Stability via Adaptive Gradient Clipping

Abstract
Loss spikes remain a persistent obstacle in large-scale language model pretraining. While previous research has attempted to identify the root cause of loss spikes by investigating individual factors, we observe that, in practice, such spikes are typically triggered by the confluence of heterogeneous factors. Empirically, loss spikes may arise from a combination of data outliers, hardware or transient computational faults, numerical precision issues, and hyperparameter settings. Regardless of the underlying cause, these spikes manifest as unstable optimizer updates, as abnormal gradients contaminate both first- and second-moment states. In this paper, we propose a principled gradient-centric remedy: AdaGC, an adaptive per-tensor gradient clipping scheme that mitigates such contamination by bounding gradient norms relative to a tensor-wise exponential moving average of their historical clipped values. AdaGC is optimizer-agnostic, introduces negligible memory overhead, and reduces communication costs compared to GlobalGC, particularly in hybrid-parallel distributed training environments. Experiments on Llama-2 7B, Mixtral 8×1B, and ERNIE 10B-A1.4B demonstrate that AdaGC robustly eliminates training instabilities, consistently reducing spike scores to zero for all models and improving downstream accuracy over GlobalGC by 1.32%, 1.27%, and 2.48%, respectively. Furthermore, AdaGC seamlessly integrates with optimizers such as Muon and Lion, consistently yielding higher average accuracy and zero spike scores.
Machine Learning, ICML
1 Introduction
(a) Mixtral
(b) ERNIE
(c) ERNIE
(d) ERNIE
Figure 1: Reproduced cases of loss spikes and mitigation via resuming. Loss spikes are triggered by (a) increasing $\beta_{2}$ or (b) reducing $\epsilon$ in AdamW, (c) using lower-precision RMSNorm, even under global gradient clipping, and (d) are resolved by resuming due to stochasticity in FlashAttention backward passes.
Table 1: Common sources of loss spikes in LLM pretraining, with typical triggers and practical mitigations.
Source Typical trigger Common mitigation
Data quality Noisy/ill-formed samples Resume and skip data batches
Hardware / transient faults GPU silent errors Fault tolerance and machine replacement
Numerical precision Activation outlier BF16 to FP32 or blockwise quantization
Optimizer / LN hyperparams AdamW/RMSNorm $\epsilon$; AdamW $\beta_{2}$ Careful hyperparameter selection
The rapid scaling of large language models (LLM) has introduced new challenges in pretraining stability, often manifesting as abrupt loss spikes or transient divergences across a wide range of model architectures and data scales (Chowdhery et al., 2023; Touvron et al., 2023; Liu et al., 2024; Team et al., 2025; Baidu-ERNIE-Team, 2025). Despite extensive empirical studies, the fundamental causes of these instabilities remain elusive. Recent research, alongside our own analyses, indicates that loss spikes can arise from a variety of sources, including: (i) data quality issues (Chowdhery et al., 2023); (ii) hardware or transient computational faults (Su, 2025); (iii) variations in numerical precision (for example, FP32 typically offers greater robustness than BF16, whereas FP8 can sometimes enhance stability by suppressing outlier values via implicit quantization (Han, 2024; Liu et al., 2024)); and (iv) the selection of optimizer and layer normalization hyperparameters, such as the $\epsilon$ parameter in RMSNorm or AdamW, and $\beta_{2}$ in AdamW (Ma et al., 2021; Cattaneo & Shigida, 2025; Bai et al., 2025). For instance, we observe that increasing $\beta_{2}$ or decreasing $\epsilon$ in AdamW can trigger loss spikes, whereas increasing the precision of RMSNorm from BF16 to FP32 significantly improves stability. In summary, Table 1 presents the common sources, typical triggers, and practical mitigations for loss spikes identified from both previous studies and our experimental findings, while Figure 1 demonstrates several representative cases we have reproduced.
Although the upstream causes of instability are diverse and often subtle, these events consistently converge at the optimizer level, manifesting as abnormal gradients. Such outlier gradients are incorporated into the optimizer’s first- and second-moment estimates, thereby corrupting parameter updates and propagating instability through subsequent training. Notably, we find that even resuming interrupted training (while keeping the random seed and data unchanged) can mitigate a loss spike, merely due to the stochastic nature of $dQ$, dKdK, and $dV$ in FlashAttention (Dao, 2023) (see Figure 1(d)). This observation further suggests that, in certain model states, even minute numerical differences can trigger a loss spike, with gradient outliers playing a critical role in both the initiation and propagation of instabilities during optimizer state updates.
While the slight stochasticity introduced by FlashAttention can sometimes circumvent a loss spike, repeatedly interrupting and resuming training imposes substantial computational overhead. Given that these instabilities stem from diverse upstream causes but ultimately converge at the optimizer level, our work does not attempt to identify the precise root causes. Instead, we adopt a gradient-centric perspective: irrespective of the initial trigger, loss spikes consistently arise when outlier gradients contaminate the optimizer states. Therefore, by preventing such gradients from entering the first- and second-moment accumulators, we provide a unified and effective strategy to mitigate training instability.
A standard mitigation strategy is global gradient clipping (GlobalGC), which bounds the global $\ell_{2}$ norm of the aggregated gradient. However, this approach is fundamentally mismatched to modern large-scale pretraining in two key respects: (1) Temporal mismatch: The optimal global clipping threshold typically decreases over the course of training; a fixed threshold risks under-clipping in later phases. (2) Spatial mismatch: Gradient statistics and rare spikes vary asynchronously across different parameter tensors, making a single global threshold insufficient—protecting one tensor may under-serve or over-constrain others.
To address these challenges, we introduce Adaptive Gradient Clipping based on Local Gradient Norm (AdaGC): a simple, per-tensor clipping rule that leverages an EMA of each tensor’s historical gradient norm as a reference. Each tensor’s gradient is clipped relative to its own EMA, preventing transient outliers from contaminating the first- and second-moment accumulators and, ultimately, the parameter updates. A brief warm-up period applies global clipping and initializes the EMA to avoid early overestimation. AdaGC is optimizer-agnostic and can be seamlessly integrated with AdamW, Lion, and Muon. Our main contributions are as follows:
  • A unified, gradient-centric perspective: We clarify how loss spikes universally propagate via abnormal gradients polluting optimizer states, irrespective of their origin, motivating intervention at the gradient level prior to moving-average accumulation.
  • An adaptive, per-tensor clipping rule: By tracking each tensor’s gradient norm statistics with an EMA, AdaGC provides both temporal adaptivity and spatial specificity, suppressing outliers while minimally disturbing typical learning dynamics.
  • System efficiency: We analyze computational and communication overhead, showing that AdaGC reduces communication relative to GlobalGC under hybrid parallel distributed training.
[…]
2 Related Work
Table 2: Comparison of major gradient/update clipping methods for stability. Here, $\boldsymbol{\theta}_{t}$ denotes the model parameters, $\boldsymbol{g}_{t}$ the gradients, $\Delta_{t}$ the optimizer update, $\eta_{t}$ the learning rate, $\lambda_{abs}$ the absolute threshold, and $\lambda_{rel}$ the relative threshold.
Method Algorithm Gradient Update Granularity Threshold Type
GlobalGC (Pascanu et al., 2013) $\min\{1.0,\lambda_{abs}\frac{1}{\|\boldsymbol{g}_{t}\|}\}$ Global Fixed constant
ClipByValue $clamp(-\lambda_{abs},\lambda_{abs})$ Element Fixed constant
AGC (Brock et al., 2021) $\min\{1.0,\lambda_{rel}\frac{\|\boldsymbol{\theta}_{t}\|}{\|\Delta_{t}\|}\}$ Unit Weight $\ell_{2}$ norm
Clippy (Tang et al., 2023) $\min\{1.0,\min(\frac{\lambda_{rel}\|\boldsymbol{\theta}_{t}\|_{\infty}+\lambda_{abs}}{\eta_{t}*\|\Delta_{t}\|_{\infty}})\}$ Tensor Weight $\ell_{\infty}$ norm
AdaGC (ours) $\min\{1.0,\lambda_{rel}\frac{\gamma_{t-1,i}}{\|\boldsymbol{g}_{t,i}\|}\}$ Tensor EMA of gradient norm
$\gamma_{t,i}=\beta\gamma_{t-1,i}+(1-\beta)\|\boldsymbol{g}_{t,i}\|$
Stability in large-scale pretraining: Dozens of approaches address instability during large-model pretraining, including: architectural advances (Pre-LN (Xiong et al., 2020), RMSNorm (Zhang & Sennrich, 2019)), careful initialization (Nguyen & Salazar, 2019; Takase et al., 2023; Nishida et al., 2024), auxiliary loss terms (Max-z loss (Yang et al., 2023)). Recent work (OLMo et al., 2024) also explores combining multiple stabilization strategies. These measures improve average stability but do not directly prevent abnormal gradients from corrupting optimizer states.
Gradient/Update Clipping: Gradient and update clipping achieve stability by limiting the magnitude of gradients and parameter updates, preventing excessively large weight updates. Global gradient clipping (Pascanu et al., 2013) is prevalent, with innovative approaches like AGC (Brock et al., 2021) and Clippy (Tang et al., 2023), which use model weights to adjust the clipping threshold. The SPAM (Huang et al., 2025) method stabilizes the model training process by introducing a momentum reset mechanism and an element-wise gradient clipping strategy based on second-moment estimation. Alternatives like Adafactor (Shazeer & Stern, 2018), StableAdamW (Wortsman et al., 2023), and LAMB (You et al., 2019) offer update clipping techniques better suited for stability training of large-scale models. Nonetheless, a significant number of loss spikes still occur during the training of large language models, even with the application of these methodologies. Due to our gradient-centric perspective, we focus our discussion on clipping-based methods. These methods fall into two categories: value-based approaches, which truncate individual gradient components exceeding a predefined limit, and norm-based approaches, which rescale the entire gradient vector only when its overall magnitude exceeds a threshold. AdaGC belongs to the norm-based category, leveraging adaptive per-tensor norm thresholds to stabilize training. For a comparative summary, see Table 2.
Figure 2: Visualization of the gradient norm, first-moment norm, second-moment norm, update norm, loss, and global gradient norm for the embedding of Llama-2 1.3B during warmup phase. Each row represents a different clipping method: the first row is GlobalGC, the second is AGC, the third is Clippy, and the fourth is our AdaGC. The black curve in each plot shows the loss trajectory.
3 Motivation: From Root-Cause Diversity to a Unified Gradient-Centric Remedy
Through a series of experiments (see Figure 1 and Figure 2), we observe that loss spikes encountered under diverse settings consistently coincide with abrupt fluctuations in the gradient norm. Comparative analyses further reveal limitations of existing methods such as GlobalGC, AGC, and Clippy: GlobalGC’s static global threshold cannot detect or suppress localized abnormal gradients, allowing outliers to contaminate optimizer states and trigger instability. AGC and Clippy focus on controlling parameter updates, leaving internal moments vulnerable to large gradient outliers.
As discussed in the Introduction (Section  1), loss spikes typically result from a combination of multiple factors. While the specific triggers may vary, these loss spikes share a common manifestation: abnormally large gradients are incorporated into the optimizer’s moment estimates, leading to unstable updates. Based on these analyses, we propose a unified remedy: regardless of the root cause, instability in large-scale training is best addressed via gradient-centric clipping. Specifically, only localized and adaptive clipping, applied before gradients are integrated into the optimizer’s moment estimates, can effectively constrain the influence of outlier gradients. We thus distill two key principles for loss spike mitigation: (1) Locality: clip gradients for each parameter tensor individually, avoiding the insensitivity of a global threshold; (2) Adaptivity: dynamically adjust each tensor’s clipping threshold, e.g., using an EMA of its recent gradient norms.
4 Methodology: AdaGC
4.1 Preliminaries
Notations. Let $x_{t}\in\mathbb{R}^{d}$ denote a parameter vector where $x_{t}^{j}$ represents its $j$-th coordinate for $j\in[d]$. We write $\nabla_{x}f(x)$ for the gradient of any differentiable function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$, and use $u^{2}$ and $u/v$ to denote element-wise square and division operations for vectors $u,v\in\mathbb{R}^{d}$. The $\ell_{2}$-norm and $\ell_{\infty}$-norm are denoted by $\|\cdot\|$ and $\|\cdot\|_{\infty}$, respectively. For asymptotic comparisons, we write $f=\mathcal{O}(g)$ if $\exists c>0$ such that $f(x)\leq cg(x)$ for all $x$ in the domain.
Gradient Clipping Fundamentals. Consider a stochastic optimization problem with parameters $\theta\in\mathbb{R}^{d}$ and loss function $f(\theta;X_{t})$ evaluated on mini-batch $X_{t}$ at step $t$. Standard gradient descent updates follow:
$\small\theta_{t}=\theta_{t-1}-\eta_{t}\nabla_{\theta}f(\theta_{t-1},X_{t})$ (1)
To prevent unstable updates from gradient explosions, GlobalGC (Pascanu et al., 2013) modifies the update rule as:
$\small\begin{array}[]{c}\theta_{t}=\theta_{t-1}-\eta_{t}h_{t}\nabla_{\theta}f(\theta_{t-1},X_{t})\\[3.6999pt] \text{ where }h_{t}:=\min\left\{\frac{\lambda_{abs}}{\|\nabla_{\theta}f(\theta_{t-1};X_{t})\|},1.0\right\}\end{array}$ (2)
Here $\lambda_{abs}$ is an absolute clipping threshold requiring careful tuning, and $\eta_{t}$ is the learning rate. Our work focuses on norm-based clipping (scaling entire gradients exceeding $\lambda_{abs}$) rather than value-based clipping (element-wise truncation).
4.2 Adaptive Gradient Clipping based on Local Gradient Norm
This section introduces a novel gradient clipping strategy termed AdaGC, which distinguishes itself by not relying on a global gradient norm. Instead, AdaGC focuses on the local gradient norm of each tensor and utilizes a dynamic adaptive mechanism for gradient clipping. The proposed method employs an EMA mechanism to maintain smoothed estimates of historical gradient norms per tensor, thus enhancing the accuracy of anomalous gradient detection and enabling independent clipping adjustments tailored to each tensor’s specific conditions. EMA is widely used in deep learning, and within AdaGC, it facilitates the balancing of historical and current gradient norms. The formulation is as follows:
$\small\begin{array}[]{c}\boldsymbol{g}_{t,i}\leftarrow h_{t,i}\cdot\boldsymbol{g}_{t,i},\text{where }h_{t,i}=\min\left\{\lambda_{rel}\frac{\gamma_{t-1,i}}{\|\boldsymbol{g}_{t,i}\|},1.0\right\},\\[9.24994pt] \gamma_{t,i}=\beta\gamma_{t-1,i}+(1-\beta)\|\boldsymbol{g}_{t,i}\|.\end{array}$ (3)
Here, $\lambda_{rel}$ is a predefined relative clipping threshold, $\boldsymbol{g}_{t,i}$ represents the gradient of the $i$-th tensor at time step $t$, and $h_{t,i}$ is a clipping function activated when $\|\boldsymbol{g}_{t,i}\|>\lambda_{rel}\cdot\gamma_{t-1,i}$, thereby scaling the gradient norm to $\lambda_{rel}\cdot\gamma_{t-1,i}$. Additionally, $\beta$ is the smoothing coefficient for EMA. We consistently incorporate the clipped gradient norm into the historical observations rather than the pre-clipped values.
Despite its simplicity, AdaGC adaptively adjusts based on the magnitude of each tensor’s gradient norm. Whenever the gradient norm at a current timestep exceeds a predefined range of average norms within a historical window, it effectively suppresses these outlier gradients.
However, during the initial stages of model training (e.g., the first 100 steps), the gradient norms are typically large and fluctuate significantly, indicating a substantial decreasing trend. Direct application of AdaGC during this period could lead to two issues: first, erroneously accumulating the early large gradient norms into the historical values, resulting in compounded errors; second, compared to GlobalGC, AdaGC might delay clipping, thus potentially slowing down the loss reduction. To address these issues, we introduce a hyperparameter $T_{start}$ (default set to 100), representing a warm-up period during which traditional GlobalGC is applied.
Additionally, AdaGC is optimizer-agnostic, can be seamlessly integrated with various optimizers, such as AdamW (Loshchilov & Hutter, 2017), Lion (Chen et al., 2024), Muon (Jordan et al., 2024), enhancing its practicality and flexibility. Algorithm 1 in Appendix A demonstrates its implementation with the AdamW optimizer.
Table 3: Zero-shot accuracy of AdaGC on Llama-2 7B under different hyperparameters.
$\lambda_{rel}$ $\beta$ 0.98 0.985 0.99 0.999
1.03 50.06 50.92 50.95 50.96
1.04 48.88 50.59 51.04 50.76
1.05 51.01 49.95 50.57 50.74
Table 4: Two-shot accuracy of AdaGC on Llama-2 7B under different hyperparameters.
$\lambda_{rel}$ $\beta$ 0.98 0.985 0.99 0.999
1.03 52.31 52.68 53.13 53.42
1.04 52.68 53.01 53.47 52.85
1.05 52.68 52.67 51.96 53.51
4.3 Memory, Computation, and Communication
Memory. As a tensor-wise method, AdaGC maintains an EMA of gradient norms for each parameter tensor, requiring storage of a single 32-bit float (4 bytes) per tensor. For ERNIE models, the total additional memory overhead has complexity of $\mathcal{O}((9+3E)\times L+3)$, where $L$ and $E$ denote the number of transformer layers and experts, respectively. Specifically, this includes four tensors from the attention module per layer, $3\times(1+E)$ tensors from the shared and router experts per layer, and two RMSNorm tensors per layer; plus one tensor each for the embedding layer, the final layer normalization, and the language modeling head. In practice, this added memory footprint is negligible compared to the overall memory requirements of large-scale model training.
Computation. The computational cost of computing $\ell_{2}$ norms is the same for both AdaGC and GlobalGC. The difference is that GlobalGC applies a uniform scaling to all gradients, while AdaGC scales each gradient tensor independently.
Communication. In setups involving data parallelism (DP), tensor parallelism (TP), and pipeline parallelism (PP), GlobalGC requires an all-reduce operation across all DP, TP, and PP groups to aggregate the global norm. In contrast, AdaGC only needs an all-reduce within each TP group to compute per-tensor local norms. This design substantially reduces communication overhead, offering increasing benefits as model and cluster sizes grow.
5 Experiments
5.1 Experimental Setup
[…]
Comparison Methods. We focus on clipping-based methods and compare gradient and update clipping baselines, including GlobalGC (Pascanu et al., 2013), Gradient Value Clipping (ClipByValue), AGC (Brock et al., 2021), and Clippy (Tang et al., 2023). We also evaluate recent methods, including SPAM (Huang et al., 2025), Scaled Embed (Takase et al., 2023), and WeSaR (Nishida et al., 2024). Results are in Appendix D.2 Table 12.
Training Details. Pre-training large-scale models is typically resource-intensive. In this study, our primary focus was to investigate training instability rather than to pursue ultimate accuracy. To facilitate multiple experimental runs, we conducted 9,000 training steps over 36 billion tokens for both Llama-2 1.3B and 7B, 36,000 steps over 36 billion tokens for Mixtral 8x1B, and 60,000 steps over 1 trillion tokens for ERNIE 10B-A1.4B, the latter additionally serving to validate the long-term stability of AdaGC. Further details regarding hyperparameters can be found in Table 9 in Appendix B.
Evaluation Metrics. To quantitatively assess training stability, we follow (OLMo et al., 2024; Karpathy, 2024) and adopt the spike score as an objective metric. Specifically, the spike score is defined as the percentage of values in a time series that deviate by at least ten standard deviations from a rolling average of the preceding 1,000 values. This metric is primarily applied to training loss to detect sudden instabilities. Additionally, we evaluate performance using the training loss and validation perplexity (PPL) curves, as well as standard benchmark results, to provide a comprehensive assessment of convergence efficiency and model quality.
Standard Benchmark. We conducted a comprehensive evaluation of the model’s zero-shot and two-shot capabilities across seven well-established benchmarks: ARC (Yadav et al., 2019), BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), OBQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2021), and MMLU (Hendrycks et al., 2020). Following standard practice (Zhang et al., 2025), we report accuracy norm for ARC-E, ARC-C, HellaSwag, OBQA, and SciQ, as well as standard accuracy for all other tasks. For ERNIE 10B-A1.4B, which has been trained on 1T tokens, we evaluate its general abilities on a range of benchmarks, including MMLU (Hendrycks et al., 2020), GSM8K (Cobbe et al., 2021), BBH (Suzgun et al., 2022), TruthfulQA (Lin et al., 2021), and HumanEval (Chen et al., 2021). These benchmarks assess the model’s enhanced capabilities in performing diverse downstream tasks, such as examination, reasoning, factuality, and coding.
(a) Llama-2 7B training dynamics.
(b) Mixtral 8x1B training dynamics.
Figure 3: Large language model training analysis: Llama-2 7B and Mixtral 8x1B model comparison shows AdaGC’s loss spike elimination and performance gains.
5.2 Critical Hyperparameter Selection
We systematically evaluated two key hyperparameters in AdaGC: the EMA coefficient $\beta$ and the relative clipping threshold $\lambda_{rel}$. Specifically, we performed a grid search on the Llama-2 7B model to optimize these two hyperparameters, using zero-shot and two-shot performance across multiple tasks as evaluation metrics. As shown in Tables 4 and 4, the best performance was achieved when $\lambda_{rel}=1.04$ and $\beta=0.99$. We therefore adopted this configuration as the default setting for subsequent experiments and terminated further hyperparameter search. In addition, as observed in Tables 4 and 4, AdaGC’s performance remains relatively stable across different hyperparameter values, suggesting that the method is robust to hyperparameter variations.
5.3 Main Experimental Results
[…]
Table 5: The Zero-Shot evaluation results of Llama-2 1.3B/7B and Mixtral 8x1B models on standard benchmarks.
Model Method ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
acc_norm acc_norm acc acc_norm acc_norm acc_norm acc acc acc_norm
Llama-2 1.3B GlobalGC 43.18 25.68 57.19 46.62 30.20 69.97 52.64 22.97 68.40 46.32
ClipByValue 42.17 25.68 59.94 44.11 30.40 69.59 53.28 22.99 68.00 46.24
Clippy 41.71 24.66 56.51 45.43 30.00 69.21 54.85 22.90 67.50 45.86
AdaGC 42.09 25.51 58.01 47.29 30.40 69.70 52.33 22.98 68.70 46.33
Llama-2 7B GlobalGC 49.49 27.56 56.30 56.06 33.60 74.59 55.33 23.12 71.20 49.69
ClipByValue 46.21 26.88 57.03 53.49 33.20 71.65 53.59 23.36 70.50 48.43
AGC 48.15 28.16 52.87 55.47 32.80 72.74 57.85 24.33 71.70 49.34
Clippy 47.69 27.73 57.46 53.34 32.40 72.74 54.38 25.36 73.40 49.39
AdaGC 49.58 28.92 57.28 57.94 32.80 74.32 58.09 23.62 76.60 51.01
Mixtral 8x1B GlobalGC 44.70 25.94 56.57 53.08 33.00 71.60 54.70 22.91 67.20 47.74
AdaGC 46.68 26.37 58.93 55.85 32.20 73.12 54.38 23.22 70.30 49.01
Table 6: Evaluation results of ERNIE 10B-A1.4B on multiple benchmarks after 21,000 (350B tokens) and 60,000 (1T tokens) training steps, comparing different optimization configurations.
Steps (tokens) Method AdamW eps MMLU GSM8K BBH TruthfulQA HumanEval Avg.
21k (350B) GlobalGC 1e-8 47.75 28.35 28.80 22.02 19.51 28.09
GlobalGC 1e-15 39.11 21.46 29.35 23.39 15.24 25.71
AdaGC 1e-15 42.07 25.32 27.89 24.92 20.73 28.19
60k (1T) GlobalGC 1e-8 48.61 39.88 30.84 30.73 22.56 34.52
GlobalGC 1e-15 48.48 40.79 30.59 28.29 23.78 34.38
AdaGC 1e-15 48.70 36.01 31.38 35.02 22.56 34.73
[…]
Table 7: Comparison of spike scores for various models under different clipping methods.
Model Llama-2 7B Mixtral 8x1B ERNIE 10B-A1.4B
Method GlobalGC ClipByValue AGC Clippy AdaGC GlobalGC AdaGC GlobalGC AdaGC
Total Steps 9K 9K 9K 9K 9K 36K 36K 21K 21K
Num Spikes 3 9 8 3 0 52 0 2 0
Spike Score (%) 0.0333 0.1000 0.0889 0.0333 0.0000 0.0144 0.0000 0.0100 0.0000
Table 8: GPU hours under the same configuration. DPS denotes distributed parallel strategies.
Model Llama-2 1.3B Llama-2 7B Mixtral 8x1B ERNIE 10B-A1.4B
DPS DP=256, TP=1, PP=1 DP=32, TP=2, PP=1 DP=256, TP=1, PP=1, EP=1 DP=64, TP=1, PP=4, EP=8
Steps 9K 9K 36K 21K
GlobalGC 513.0 1468.2 2060.8 22922
AdaGC 511.6 1402.4 2035.2 22572
[…]
5.4 Optimizer Compatibility: Muon and Lion
AdaGC is an optimizer-agnostic gradient clipping method that can be seamlessly integrated not only with AdamW, but also with other optimizers. To verify the generality of AdaGC, we conducted experiments on both LLM and VLM tasks by combining Llama-2 1.3B and CLIP ViT-Base models with the Muon and Lion optimizers, respectively, and compared them against GlobalGC. Although no loss spikes were observed under either experimental setting, AdaGC consistently demonstrated strong compatibility and generalization. In downstream zero-shot average accuracy, AdaGC outperformed GlobalGC by 0.14% (47.18% vs. 47.04%) with Muon and by 0.16% (40.81% vs. 40.65%) with Lion. These results further confirm that AdaGC can be effectively applied across different optimizers, providing stable training and improved downstream performance.
5.5 End-to-End training Wall-clock
Table 8 compares the GPU hours required for training various models using different distributed parallelism strategies. Compared to GlobalGC, AdaGC reduces end-to-end GPU hours by 0.27% on Llama-2 1.3B, 4.48% on Llama-2 7B, 1.24% on Mixtral 8x1B, and 1.53% on ERNIE 10B-A1.4B, thanks to reduced communication overhead. This highlights AdaGC’s additional communication and efficiency benefits in large-scale distributed training.
5.6 Ablation Study
We conduct systematic ablation studies across three critical dimensions of AdaGC: (1) EMA gradient norm initialization strategies, (2) GlobalGC warm-up steps, (3) adaptivity efficacy, and (4) locality granularity.
(a) Training dynamics.
(b) Average accuracy.
Figure 4: AdaGC with Muon on Llama-2 1.3B.
(c) Training dynamics.
(d) Average accuracy.
Figure 5: AdaGC with Lion on CLIP ViT-Base.
(a) $\gamma_{t,i}$ initialization.
(b) $T_{start}$ warm-up.
(c) Adaptivity, locality.
Figure 6: Training dynamics of AdaGC ablations, showing (a) the influence of different EMA initialization strategies; (b) the impact of the GlobalGC warm-up steps $T_{start}$; and (c) the effects of adaptivity and locality granularity on gradient clipping efficacy and final loss.
EMA Initialization Strategy. The initialization of EMA gradient norms requires careful design due to large initial gradient fluctuations during early training phases (first 100 steps). We evaluate five initialization variants: The default AdaGC strategy employs GlobalGC during warm-up while tracking minimum per-parameter norms ($\gamma_{t,i}=\min(\|\boldsymbol{g}_{t,i}\|,\gamma_{t-1,i})$). Comparative approaches include: (1) norm initialization without GlobalGC warm-up (directly using $\gamma_{t,i}=\min(\|\boldsymbol{g}_{t,i}\|,\gamma_{t-1,i})$ from step 0), (2) constant initialization ($\gamma_{0,i}\in\{0.5,1.0\}$), and (3) thresholded initialization ($\gamma_{t,i}=\min(\|\boldsymbol{g}_{t,i}\|,0.1)$). Figure 6(a) demonstrates that while all variants eliminate loss spikes, convergence quality varies within 0.36%. The default strategy achieves optimal final loss (2.9708 vs 2.9725 for next-best), showing that GlobalGC-guided warm-up better preserves parameter update consistency than direct initialization. This establishes the importance of phased initialization for gradient norm adaptation.
Warm-up Steps $T_{start}$. To further investigate whether the choice of GlobalGC warm-up steps $T_{start}$ has a significant impact and to provide practical guidance for practitioners, we additionally evaluate $T_{start}=\{0,50,100,150,200,500,1000,2000\}$. The results in Figure 6(b) show that $T_{start}=100$ consistently achieves the best performance. According to the EMA initialization formula $\gamma_{t,i}=\min(\|\boldsymbol{g}_{t,i}\|,\gamma_{t-1,i})$, an excessively large $T_{start}$ accumulates lower $\gamma_{t,i}$ values due to early training dynamics, which may lead to over-clipping and suppressed convergence in later training. Conversely, an overly small $T_{start}$ accumulates larger $\gamma_{t,i}$ values, which may delay clipping and hinder timely suppression of abnormal gradients. In contrast, $T_{start}=100$ introduces negligible additional overhead for large-scale training while providing consistently stable performance improvements.
Adaptivity Efficacy and Locality Granularity. We conduct three sets of ablation experiments to evaluate the adaptivity and locality of AdaGC. The baseline uses GlobalGC (no adaptivity, no locality) with a fixed threshold of 1.0. In comparison, we examine (1) adaptive global gradient norm clipping (Global AdaGC, adaptive but non-local), which employs a single adaptive threshold for the entire model, (2) tensor-wise gradient norm clipping (TensorWiseGC, local but non-adaptive), which allocated each tensor’s fixed clipping threshold proportionally to its parameter count relative to the entire model, and (3) tensor-wise adaptation (AdaGC, adaptive and local), which adjusts thresholds independently for each tensor. As shown in Figure 6(c), Global AdaGC reduces but does not completely eliminate spike events (1 event vs. 0 for tensor-wise) and yields a 0.25% higher final loss (2.9639 vs. 2.9712). Although TensorWiseGC also mitigates loss spikes, it noticeably slows down convergence and requires careful per-tensor threshold tuning to perform well. These results demonstrate that tensor-wise adaptive clipping provides both greater spike suppression and lower loss than other approaches.
6 Conclusion
The factors triggering loss spikes in large-scale pretraining are diverse and remain an open research problem, with no unified solution to date. Unlike prior work that seeks to identify root causes, we focus on a gradient-centric remedy and introduce AdaGC, an adaptive per-tensor gradient clipping method that prevents abnormal gradients from contaminating optimizer states. This approach ensures smoother updates and effectively eliminates loss spikes. Extensive experiments demonstrate that AdaGC delivers robust and stable training across both dense and MoE models, from 1.3B to 10B parameters, consistently reducing spike scores to zero and improving benchmark performance. Our results highlight AdaGC as a simple and effective solution for stable large-scale model pretraining.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix Contents
(Run LaTeX twice to generate Appendix Contents.)
Appendix A Pseudocode of AdamW with AdaGC
Algorithm 1 presents the pseudocode of AdamW integrated with AdaGC. For clearer exposition, we highlight different components according to their origins: orange indicates the procedures inherited from the original GlobalGC algorithm, while blue is used to denote the new contributions and modifications introduced by AdaGC. Specifically, the GlobalGC steps include the global gradient clipping implemented via the scaling factor and the use of the clipped gradient in subsequent moments. The AdaGC components mainly comprise adaptive per-parameter clipping, the initialization and update of the adaptive threshold $\gamma_{t,i}$, and the warm-up strategy governed by $T_{start}$.
Algorithm 1 AdamW with AdaGC
1:  given: $\{\eta_{t}\}_{t=1}^{T},\lambda_{w},\epsilon,\beta_{1},\beta_{2},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\beta\in[0,1)},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\lambda_{abs}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}T_{start}}$
2:  initialize: $\boldsymbol{\theta}_{0},m_{0}\leftarrow 0,v_{0}\leftarrow 0,t\leftarrow 0$
3:  repeat
4:  compute $\boldsymbol{g}_{t}$ = $\nabla_{\boldsymbol{\theta}}f_{t}(\boldsymbol{\theta}_{t-1},X_{t})$
5:  if $t<T_{start}$ then
6:   $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}h_{t}=\min\left\{\frac{\lambda_{abs}}{\|\boldsymbol{g}_{t}\|},1.0\right\}$}
7:   $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\widehat{\boldsymbol{g}_{t}}=h_{t}\cdot\boldsymbol{g}_{t}$}
8:   for $i\in|\boldsymbol{\theta}|$ do
9:    $^{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma_{t,i}=\min\left\{\gamma_{t-1,i},\|\widehat{\boldsymbol{g}_{t,i}}\|\right\},\gamma_{0,i}=\|\widehat{\boldsymbol{g}_{1,i}}\|}$
10:   end for
11:  else
12:   for $i\in|\boldsymbol{\theta}|$ do
13:    $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}h_{t,i}=\min\left\{\lambda_{rel}\frac{\gamma_{t-1,i}}{\|\boldsymbol{g}_{t,i}\|},1.0\right\}$}
14:    $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\widehat{\boldsymbol{g}_{t,i}}=h_{t,i}\cdot\boldsymbol{g}_{t,i}$}
15:    $^{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\gamma_{t,i}=\beta\gamma_{t-1,i}+(1-\beta)\|\widehat{\boldsymbol{g}_{t,i}}\|}$
16:   end for
17:  end if
18:  $\boldsymbol{m}_{t}=\beta_{1}\boldsymbol{m}_{t-1}+(1-\beta_{1}){\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\widehat{\boldsymbol{g}_{t}}}$
19:  $\boldsymbol{v}_{t}=\beta_{2}\boldsymbol{v}_{t-1}+(1-\beta_{2}){\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\widehat{\boldsymbol{g}_{t}}^{2}}$
20:  $\widehat{\boldsymbol{m}_{t}}=\boldsymbol{m}_{t}/(1-\beta_{1}^{t}),\quad\widehat{\boldsymbol{v}_{t}}=\boldsymbol{v}_{t}/(1-\beta_{2}^{t})$
21:  $\boldsymbol{\theta}_{t}=\boldsymbol{\theta}_{t-1}-\eta_{t}\lambda_{w}\boldsymbol{\theta}_{t-1}-\eta_{t}\widehat{\boldsymbol{m}_{t}}/(\sqrt{\widehat{\boldsymbol{v}_{t}}}+\epsilon)$
22:  until $\boldsymbol{\theta}_{t}$ not converge
Appendix B Hyper-Parameters
B.1 Model Hyper-Parameters
Table 9 summarizes the model hyper-parameters used for all experiments. For each model, we report the core architecture settings (such as number of layers, hidden dimension, attention heads, and feedforward dimension), MoE-related configurations, and main optimization hyper-parameters (including learning rate, warmup, weight decay, and Adam parameters). Clipping thresholds $\lambda_{abs}$, $\lambda_{rel}$, and momentum $\beta$ are also listed, in correspondence with the techniques discussed in the main text. All experiments use a batch size and sequence length as shown, and we employ bfloat16 precision for most models except ERNIE, which uses float8. The symbol ‘–’ indicates settings not applicable to a specific architecture.
Table 9: Hyper-parameters used in our LLMs experiments. $\lambda_{abs}$ represents the absolute global clipping threshold of GlobalGC. $\lambda_{rel}$ and $\beta$ represent the relative clipping threshold and the momentum of our AdaGC, respectively. The symbol ‘–’ indicates that the parameter is not applicable.
Model Llama-2 1.3B Llama-2 7B ERNIE 10B-A1.4B Mixtral 8x1B
Precision bfloat16 bfloat16 float8 bfloat16
Num layers 24 32 25 24
Hidden dim size 2048 4096 2560 2048
FFN dim size 5461 11008 1024 5632
Num attention heads 32 32 20 32
Num key value heads 32 32 4 4
Attention bias
Num shared experts - - 1 0
Num router experts - - 48 8
Num experts per token - - 3 2
Sequence length 2048 2048 4096 2048
Batch size 2048 2048 4096 512
Iterations 9000 9000 21000 36000
Learning rate $3.0\times 10^{-4}$ $3.0\times 10^{-4}$ $3.0\times 10^{-4}$ $3.0\times 10^{-4}$
LR decay cosine cosine wsd cosine
Warmup iterations 2000 2000 2000 500
Weight decay 0.1 0.1 0.1 0.1
Adam $\beta_{1}$ 0.90 0.90 0.90 0.90
Adam $\beta_{2}$ 0.95 0.95 0.95 0.999
$\lambda_{abs}$ 1.0 1.0 1.0 1.0
$\lambda_{rel}$ 1.04 1.04 1.04 1.04
$\beta$ 0.99 0.99 0.99 0.99
B.2 Clipping Hyper-Parameters
For other clipping methods, we primarily followed the recommended default settings from prior work, and performed limited tuning only when necessary to ensure a fair comparison.
Specifically:
  • GlobalGC: We used the commonly adopted global clipping threshold $\lambda_{abs}=1.0$ in large-scale pretraining.
  • ClipByValue: Following the SPAM (Huang et al., 2025) setting, we set the clipping threshold to $\lambda_{abs}=1e-3$.
  • AGC: We performed small-range tuning over $\lambda_{rel}\in\{1e-2,1e-3,1e-4\}$ to find the best setting.
  • Clippy: We tuned over $\lambda_{abs}\in\{0.1,0.3,0.5\}$ and $\lambda_{rel}\in\{1e-2,1e-3,1e-4\}$ to select the optimal combination.
  • SPAM: We adopted the default hyperparameters recommended for standard pretraining in the original paper, which were reported to perform well across diverse settings. Specifically, we set the interval to $\Delta T=500$, the threshold to $\theta=5000$, and the warmup steps to $N=150$.
The final hyper-parameters used for other clipping methods are summarized in Table 10.
Table 10: Hyper-parameters for other clipping methods.
Method Hyperparameters
GlobalGC $\lambda_{abs}=1.0$
ClipByValue $\lambda_{abs}=1e-3$
AGC $\lambda_{rel}=1e-3$
Clippy $\lambda_{abs}=0.5,\lambda_{rel}=1e-2$
SPAM $\Delta T=500$, $\theta=5000$, $N=150$
Appendix C Experimental Details for CLIP
To further investigate the optimizer compatibility of AdaGC, we evaluated its effect on large-scale vision-language model pre-training, focusing on the CLIP ViT-Base model (Radford et al., 2021) with the Lion optimizer (Chen et al., 2024). The model comprises 151 million parameters and is trained on the LAION-400M (Schuhmann et al., 2021) dataset. Training is conducted for 20,000 steps, covering 320M image-text pairs.
The key training hyper-parameters are as follows: a learning rate of 0.002, weight decay of 0.2, and batch size of 32,768. We employ patch-dropout with a drop rate of 0.5 (Li et al., 2023), following recent best practices (Wortsman et al., 2023). The learning rate is linearly warmed up for the first 5,000 steps (Goyal et al., 2017), and subsequently decayed according to a cosine schedule (Loshchilov & Hutter, 2016).
Following pre-training, we report downstream zero-shot evaluation results on the ImageNet (Russakovsky et al., 2015) validation set. The results are shown in Figure 5 in the main text.
Appendix D More Evaluation Results
D.1 Results on Downstream Benchmarks
The Two-Shot evaluation results of Llama-2 1.3B/7B and Mixtral 8x1B models on standard benchmarks are presented in Table 11.
Table 11: The Two-Shot evaluation results of Llama-2 1.3B/7B and Mixtral 8x1B models on standard benchmarks. The best scores in each column are bolded. HellaSw. = HellaSwag, W.G. = WinoGrande.
Model Method ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
acc_norm acc_norm acc acc_norm acc_norm acc_norm acc acc acc_norm
Llama-2 1.3B GlobalGC 47.26 25.60 50.31 46.44 32.20 69.64 52.33 25.07 77.80 47.41
ClipByValue 47.10 25.77 56.54 43.97 30.00 68.88 52.96 26.09 77.20 47.61
Clippy 46.55 25.85 49.76 45.71 30.00 70.02 53.20 25.69 77.70 47.16
AdaGC 46.04 26.19 49.72 47.51 31.00 69.70 54.38 24.98 78.50 47.56
Llama-2 7B GlobalGC 55.81 28.58 60.70 56.54 33.00 73.72 56.75 25.51 83.20 52.64
ClipByValue 51.94 26.88 57.55 53.36 32.40 72.31 54.14 26.63 81.60 50.75
AGC 52.95 28.67 56.15 55.69 35.40 73.07 56.43 26.88 82.80 52.00
Clippy 52.86 29.10 56.48 53.76 31.80 73.07 55.72 26.03 82.60 51.27
AdaGC 56.86 29.61 59.36 57.89 33.60 73.99 57.62 26.46 85.90 53.47
Mixtral 8x1B GlobalGC 50.34 27.39 58.81 52.96 34.20 71.16 54.06 25.37 79.90 50.47
AdaGC 53.83 28.42 58.69 55.66 33.80 73.07 54.14 25.12 81.80 51.61
D.2 Results of Other Baseline Methods
Table 12: The Zero-Shot evaluation results of Llama-2 1.3B/7B models on standard benchmarks.
Model Method ARC-E ARC-C BoolQ HellaSw. OBQA PIQA W.G. MMLU SciQ Avg.
acc_norm acc_norm acc acc_norm acc_norm acc_norm acc acc acc_norm
Llama-2 1.3B WeSaR-GlobalGC 43.56 25.17 59.94 45.08 30.00 70.29 52.96 22.90 65.80 46.19
SPAM 42.05 24.83 59.60 42.82 30.00 69.31 52.17 23.02 66.40 45.58
ScaledEmbed-GlobalGC 42.21 25.51 59.66 45.50 31.80 70.02 53.28 23.22 65.20 46.27
AdaGC 42.09 25.51 58.01 47.29 30.40 69.70 52.33 22.98 68.70 46.33
Llama-2 7B WeSaR-GlobalGC 49.75 27.22 56.12 55.38 33.80 73.39 56.27 23.02 71.40 49.59
SPAM 48.53 25.77 60.34 51.89 32.60 72.03 54.54 22.95 71.00 48.85
ScaledEmbed-GlobalGC 48.57 26.71 60.89 54.32 32.60 72.25 55.33 23.66 70.50 49.42
AdaGC 49.58 28.92 57.28 57.94 32.80 74.32 58.09 23.62 76.60 51.01
Table 13: Comparison of spike scores for various models under different methods.
Model Llama-2 1.3B Llama-2 7B
Method WeSaR-GlobalGC SPAM ScaledEmbed-GlobalGC AdaGC WeSaR-GlobalGC SPAM ScaledEmbed-GlobalGC AdaGC
Total Steps 9K 9K 9K 9k 9K 9K 9K 9K
Num Spikes 2 0 10 0 1 3 8 0
Spike Score (%) 0.0222 0.0000 0.1111 0.0000 0.0111 0.0333 0.0889 0.0000
In addition to the clipping-based baselines discussed in the main text, we also compare AdaGC with several recent methods that aim to improve the stability and generalization of large language model (LLM) training, including SPAM (Huang et al., 2025), Scaled Embed (Takase et al., 2023), and WeSaR (Nishida et al., 2024). The detailed results under the zero-shot setting and spike score are summarized in Table 12 and 13. The training dynamics are shown in Figures 8 and 9.
Among these methods, SPAM is designed to stabilize training by adjusting the optimizer’s behavior, while Scaled Embed and WeSaR focus on initialization or embedding scaling strategies to suppress loss spikes. Our experiments show that, although some of these methods can partly mitigate instability or improve certain metrics, AdaGC generally achieves higher stability and better final performance across model scales. Notably, while WeSaR is also effective at suppressing loss spikes, its reliance on special parameter initialization limits its applicability to from-scratch training. In contrast, AdaGC works reliably under both from-scratch and resumed training regimes, providing stronger flexibility. Overall, these results demonstrate AdaGC’s superior robustness and generalization compared to other non-clipping baselines.
Appendix E More Visualization Results
E.1 Training Dynamics
Figure 7: Llama-2 1.3B training dynamics of clipping methods.
Figure 8: Llama-2 1.3B training dynamics of other baseline methods.
Figure 9: Llama-2 7B training dynamics of other baseline methods.
E.2 Optimizer State Dynamics
Figure 10: Visualization of the gradient norm, first-moment norm, second-moment norm, update norm, loss, and global gradient norm for the embedding of Llama-2 1.3B. Each row represents a different clipping method: the first row is GlobalGC, the second is AGC, the third is Clippy, and the fourth is our AdaGC. The black curve in each plot shows the loss trajectory.
Figure 11: Visualization of the gradient norm, first-moment norm, second-moment norm, update norm, loss, and global gradient norm for the encoder_layers_3_self_attention_query_key_value of Llama-2 1.3B. Each row represents a different clipping method: the first row is GlobalGC, the second is AGC, the third is Clippy, and the fourth is our AdaGC. The black curve in each plot shows the loss trajectory.
Figure 12: Visualization of the gradient norm, first-moment norm, second-moment norm, update norm, loss, and global gradient norm for the LMHead of Llama-2 1.3B. Each row represents a different clipping method: the first row is GlobalGC, the second is AGC, the third is Clippy, and the fourth is our AdaGC. The black curve in each plot shows the loss trajectory.
Figure 13: Visualization of the gradient norm, first-moment norm, second-moment norm, update norm, loss, and global gradient norm for the encoder_final_layernorm of Llama-2 1.3B. Each row represents a different clipping method: the first row is GlobalGC, the second is AGC, the third is Clippy, and the fourth is our AdaGC. The black curve in each plot shows the loss trajectory.
Figure 14: Visualization of the gradient norm, first-moment norm, second-moment norm, update norm, loss, and global gradient norm for the encoder_layers_0_self_attention_query_key_value of Llama-2 1.3B. Each row represents a different clipping method: the first row is GlobalGC, the second is AGC, the third is Clippy, and the fourth is our AdaGC. The black curve in each plot shows the loss trajectory.
Figure 15: Visualization of the gradient norm, first-moment norm, second-moment norm, update norm, loss, and global gradient norm for the encoder_layers_23_input_layernorm of Llama-2 1.3B. Each row represents a different clipping method: the first row is GlobalGC, the second is AGC, the third is Clippy, and the fourth is our AdaGC. The black curve in each plot shows the loss trajectory.
Appendix F Convergence Analysis and Proof
Any operation that modifies gradients may potentially result in non-convergence. In this section, rather than providing a theoretical guarantee that AdaGC eliminates loss spikes, we present the convergence guarantee for Adam with AdaGC, stated as follows:
Theorem F.1.
Under mild assumptions, by selecting $\alpha_{t}=\mathcal{O}(1/\sqrt{T})$, $\beta_{2}=1-\mathcal{O}(1/T)$ and $\beta_{1}<\sqrt{\beta_{2}}$, when $\tau$ is randomly chosen from $\{1,2,\cdots,T\}$ with equal probabilities, it holds that
$\mathbb{E}\|\nabla f(\theta_{\tau})\|^{2}=\mathcal{O}\left(\frac{1}{\sqrt{T}}\right).$
Theorem F.1 shows that even with local clipped gradient, Adam with AdaGC can converge at the same rate as vanilla Adam (Kingma & Ba, 2014).
In the following, we provide the necessary assumptions and lemmas for the proofs of Theorem F.1.
Notations
The $k$-th component of a vector $v_{t}$ is denoted as ${v}_{t,k}$. Other than that, all computations that involve vectors shall be understood in the component-wise way. We say a vector $v_{t}\geq 0$ if every component of $v_{t}$ is non-negative, and $v_{t}\geq w_{t}$ if $v_{t,k}\geq w_{t,k}$ for all $k=1,2,\ldots,d$. The $\ell_{1}$ norm of a vector $v_{t}$ is defined as $\|v_{t}\|_{1}=\sum_{k=1}^{d}|{v}_{t,k}|$. The $\ell_{2}$ norm is defined as $\|v_{t}\|^{2}=\langle v_{t},v_{t}\rangle=\sum_{k=1}^{d}|{v}_{t,k}|^{2}$. Given a positive vector $\hat{\eta}_{t}$, it will be helpful to define the following weighted norm: $\|v_{t}\|^{2}_{\eta_{t}}=\langle v_{t},\hat{\eta}_{t}v_{t}\rangle=\sum_{k=1}^{d}\hat{\eta}_{t,k}|{v}_{t,k}|^{2}$.
Assumption F.2.
The function f is lower bounded by $\underline{f}$ with L-Lipschitz gradient.
Assumption F.3.
The gradient estimator $g$ is unbiased with bounded norm, e.g,
$\begin{split}&\mathbb{E}[g|x_{t}]=\nabla f(x_{t}),\ \|g_{t}\|\leq G.\end{split}$
Assumption F.4.
The coefficient of clipping $h_{t,i}$ is lower bounded by some $h_{0}>0$.
Assumption F.5.
$\|g_{t}-\nabla f(x_{t})\|\leq p\|\nabla f(x_{t})\|$ holds for some $p<1$ and for all t.
Remark F.6.
Assumption F.2 and Assumption F.3 are widely used in the proof of optimization algorithm with adaptive learning rates (Reddi et al., 2018). Assumption F.4 is because the gradient norm changes slowly when training the neural network, and the last assumption holds when the batch size is large enough.
Lemma F.7.
Let $\zeta:=\beta_{1}^{2}/{\beta_{2}}$. We have the following estimate
$m_{t}^{2}\leq\frac{1}{(1-\zeta)(1-\beta_{2})}v_{t},~\forall t.$ (4)
Proof.
By the iteration formula $m_{t}=\beta_{1}m_{t-1}+(1-\beta_{1})\hat{g}_{t}$ and $m_{0}=0$, we have
$m=\sum_{i=1}^{t}\beta_{1}^{t-i}(1-\beta_{1})\hat{g}_{i}.$
Similarly, by $v_{t}=\beta_{2}v_{t-1}+(1-\beta_{2})\hat{g}_{t}^{2}$ and $v_{0}=0$, we have
$v_{t}=\sum_{i=1}^{t}\beta_{2}^{t-i}(1-\beta_{2})\hat{g}_{i}^{2}$
It follows by arithmetic inequality that
$\begin{split}m_{t}^{2}&=\left(\sum_{i=1}^{t}\frac{(1-\beta_{1})\beta_{1}^{t-i}}{\sqrt{(1-\beta_{2})\beta_{2}^{t-i}}}\sqrt{(1-\beta_{2})\beta_{2}^{t-i}}\hat{g}_{i}\right)^{2}\\ &\leq\left(\sum_{i=1}^{t}\frac{(1-\beta_{1})^{2}\beta_{1}^{2(t-i)}}{(1-\beta_{2})\beta_{2}^{t-i}}\right)\left(\sum_{i=1}^{t}(1-\beta_{2})\beta_{2}^{t-i}\hat{g}_{i}^{2}\right)=\left(\sum_{i=1}^{t}\frac{(1-\beta_{1})^{2}\beta_{1}^{2(t-i)}}{(1-\beta_{2})\beta_{2}^{t-i}}\right)v_{t}.\end{split}$
Further, we have
$\sum_{i=1}^{t}\frac{(1-\beta_{1})^{2}\beta_{1}^{2(t-i)}}{(1-\beta_{2})\beta_{2}^{t-i}}\leq\frac{1}{1-\beta_{2}}\sum_{i=1}^{t}\left(\frac{\beta_{1}^{2}}{\beta_{2}}\right)^{t-i}=\frac{1}{1-\beta_{2}}\sum_{k=0}^{t-1}\zeta^{k}\leq\frac{1}{(1-\zeta)(1-\beta_{2})}.$
The proof is completed. ∎
Lemma F.8.
The following estimate holds
$\sum_{t=1}^{T}\|\Delta_{t}\|^{2}\leq\frac{\alpha^{2}G^{2}}{\epsilon}$
Proof.
By using the definition of $m_{t}$, it holds $\|m_{t}\|^{2}\leq G^{2}$.
Then, $\|\Delta_{t}\|^{2}=\|\frac{\alpha_{t}m_{t}}{\sqrt{v_{t}}+\epsilon}\|^{2}\leq\frac{G^{2}}{\epsilon}\alpha_{t}^{2}$ by using the definition of $\Delta_{t}$.
Therefore, $\sum_{t=1}^{T}\|\Delta_{t}\|^{2}\leq\frac{G^{2}}{\epsilon}\sum_{t=1}^{T}\frac{\alpha^{2}}{T}=\frac{G^{2}\alpha^{2}}{\epsilon}$.
Lemma F.9.
With the Assumption F.4 and F.5, it holds that
$\mathbb{E}\left\langle\nabla f\left(\theta_{t}\right),\hat{\eta}_{t}\hat{g}_{t}\right\rangle\geq h_{0}\mathbb{E}\left\|\nabla f\left(\theta_{t}\right)\right\|_{\hat{\eta}_{t}}^{2}.$
Proof.
According to Assumption F.5, it holds that
$\begin{split}\left\langle\nabla_{i}f\left(\theta_{t}\right),g_{t,i}\right\rangle&=-\frac{1}{2}\left(\left\|\nabla_{i}f\left(\theta_{t}\right)-g_{t,i}\right\|^{2}-\left\|\nabla_{i}f\left(\theta_{t}\right)\right\|^{2}-\left\|g_{t,i}\right\|^{2}\right)\\ &\geq(1-p^{2})\left\|\nabla_{i}f\left(\theta_{t}\right)\right\|^{2}\geq 0.\end{split}$
Thus, it holds that
$\begin{split}\mathbb{E}\left[\left\langle\nabla f(x_{t}),\hat{\eta}_{t}\hat{g}_{t}\right\rangle\right]&=\mathbb{E}\left[\sum_{i}\left\langle\nabla_{i}f(\theta_{t}),h_{t,i}\hat{\eta}_{t,i}g_{t,i}\right\rangle\right]\\ &\geq h_{0}\mathbb{E}\left[\sum_{i}\left\langle\nabla_{i}f(x_{t}),h_{t,i}\hat{\eta}_{t,i}g_{t,i}\right\rangle\right]\\ &=h_{0}\mathbb{E}\left\langle\nabla f(\theta_{t}),\hat{\eta}_{t}g_{t}\right\rangle=h_{0}\mathbb{E}\|\nabla f(\theta_{t})\|^{2}_{\hat{\eta}_{t}}.\end{split}$
Let $\Delta_{t}:=\theta_{t+1}-\theta_{t}=-\alpha_{t}m_{t}/(\sqrt{v_{t}}+\epsilon)$. Let $\hat{v}_{t}=\beta_{2}v_{t-1}+(1-\beta_{2})\delta_{t}^{2}$, where $\delta_{t}^{2}=\mathbb{E}_{t}\left[\hat{g}_{t}^{2}\right]$ and let $\hat{\eta}_{t}=\alpha_{t}/\sqrt{\hat{v}_{t}+\epsilon}$.
Lemma F.10.
Let $M_{t}=\mathbb{E}\left[\langle\nabla f(\theta_{t}),\Delta_{t}\rangle+L\|\Delta_{t}\|^{2}\right]$. Let $\alpha_{t}=\alpha/\sqrt{T}$ and $\beta_{2}=1-\beta/T$. Then, for $T\geq 1$ we have
$\begin{split}\sum_{t=1}^{T}M_{t}\leq\frac{C_{2}}{1-\sqrt{\zeta}}+\frac{LG^{2}\alpha^{2}}{(1-\sqrt{\zeta})\epsilon}-\frac{(1-\beta_{1})h_{0}}{2}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\theta_{t})\|_{\hat{\eta}_{t}}^{2},\end{split}$ (5)
where $C_{2}=\frac{5}{2(1-\beta_{1})h_{0}}\left((1-\beta_{1})^{2}\frac{4\alpha\beta G^{4}}{\epsilon^{3}}+\beta_{1}^{2}\alpha\beta\left(\frac{G^{4}}{\beta_{2}\epsilon^{3}}+\frac{(1+\epsilon)G^{2}}{(1-\zeta)\epsilon\beta_{2}}+\frac{G^{4}}{\beta_{2}}\right)\right).$
Proof.
To split $M_{t}$, firstly we introduce the following two equalities. Using the definitions of $v_{t}$ and $\hat{v}_{t}$, we obtain
$\begin{split}\frac{\left(1-\beta_{1}\right)\alpha_{t}\hat{g}_{t}}{\sqrt{v_{t}}+\epsilon}&=\frac{\left(1-\beta_{1}\right)\alpha_{t}\hat{g}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon}+\left(1-\beta_{1}\right)\alpha_{t}\hat{g}_{t}\left(\frac{1}{\sqrt{v_{t}}+\epsilon}-\frac{1}{\sqrt{\hat{v}_{t}}+\epsilon}\right)\\ &=\left(1-\beta_{1}\right)\hat{\eta}_{t}\hat{g}_{t}+\left(1-\beta_{1}\right)\alpha_{t}\hat{g}_{t}\frac{\left(1-\beta_{2}\right)\left(\sigma_{t}^{2}-\hat{g}_{t}^{2}\right)}{\left(\sqrt{v_{t}}+\epsilon\right)\left(\sqrt{\hat{v}_{t}}+\epsilon\right)\left(\sqrt{v_{t}}+\sqrt{\hat{v}_{t}}\right)}\\ &=\left(1-\beta_{1}\right)\hat{\eta}_{t}\hat{g}_{t}+\left(1-\beta_{1}\right)\hat{\eta}_{t}\hat{g}_{t}\frac{\left(1-\beta_{2}\right)\left(\sigma_{t}^{2}-\hat{g}_{t}^{2}\right)}{\left(\sqrt{v_{t}}+\epsilon\right)\left(\sqrt{v_{t}}+\sqrt{\hat{v}_{t}}\right)}\end{split}$
In addition, we can obtain:
$\begin{split}&\beta_{1}\alpha_{t}m_{t-1}\left(\frac{1}{\sqrt{\beta_{2}v_{t-1}}+\sqrt{\beta_{2}}\epsilon}-\frac{1}{\sqrt{v_{t}}+\epsilon}\right)\\ &=\beta_{1}\alpha_{t}m_{t-1}\frac{\left(1-\beta_{2}\right)\hat{g}_{t}^{2}}{\left(\sqrt{v_{t}}+\epsilon\right)\left(\sqrt{\beta_{2}v_{t-1}}+\sqrt{\beta_{2}}\epsilon\right)\left(\sqrt{v_{t}}+\sqrt{\beta_{2}v_{t-1}}\right)}+\beta_{1}\alpha_{t}m_{t-1}\frac{\left(1-\sqrt{\beta_{2}}\right)\epsilon}{\left(\sqrt{v_{t}}+\epsilon\right)\left(\sqrt{\beta_{2}v_{t-1}}+\sqrt{\beta_{2}}\epsilon\right)}\\ &=\beta_{1}\alpha_{t}m_{t-1}\frac{\left(1-\beta_{2}\right)\hat{g}_{t}^{2}}{\left(\sqrt{\hat{v}_{t}}+\epsilon\right)\left(\sqrt{\beta_{2}v_{t-1}}+\sqrt{\beta_{2}}\epsilon\right)\left(\sqrt{v_{t}}+\sqrt{\beta_{2}v_{t-1}}\right)}\\ &\qquad+\beta_{1}\alpha_{t}m_{t-1}\frac{\left(1-\beta_{2}\right)\hat{g}_{t}^{2}}{\left(\sqrt{\beta_{2}v_{t-1}}+\sqrt{\beta_{2}}\epsilon\right)\left(\sqrt{v_{t}}+\sqrt{\beta_{2}v_{t-1}}\right)}\left(\frac{1}{\sqrt{\hat{v}_{t}}+\epsilon}-\frac{1}{\sqrt{v_{t}}+\epsilon}\right)\\ &\qquad+\beta_{1}\alpha_{t}m_{t-1}\frac{\left(1-\sqrt{\beta_{2}}\right)\epsilon}{\left(\sqrt{\hat{v}_{t}}+\epsilon\right)\left(\sqrt{\beta_{2}v_{t-1}}+\sqrt{\beta_{2}}\epsilon\right)}+\beta_{1}\alpha_{t}m_{t-1}\frac{\left(1-\sqrt{\beta_{2}}\right)\epsilon}{\sqrt{\beta_{2}v_{t-1}}+\sqrt{\beta_{2}}\epsilon}\left(\frac{1}{\sqrt{\hat{v}_{t}}+\epsilon}-\frac{1}{\sqrt{v_{t}}+\epsilon}\right)\\ &=\beta_{1}m_{t-1}\hat{\eta}_{t}\frac{\left(1-\beta_{2}\right)\hat{g}_{t}^{2}}{\left(\sqrt{\beta_{2}v_{t-1}}+\sqrt{\beta_{2}}\epsilon\right)\left(\sqrt{v_{t}}+\sqrt{\beta_{2}v_{t-1}}\right)}\\ &\qquad+\beta_{1}\hat{\eta}_{t}m_{t-1}\frac{(1-\beta_{2})^{2}\hat{g}_{t}^{2}(\sigma_{t}^{2}-\hat{g}_{t}^{2})}{(\sqrt{v_{t}}+\epsilon)(\sqrt{v_{t}}+\sqrt{\hat{v}_{t}})(\sqrt{\beta_{2}v_{t-1}}+\sqrt{\beta_{2}}\epsilon)(\sqrt{v_{t}}+\sqrt{\beta_{2}v_{t-1}})}\\ &\qquad+\beta_{1}\hat{\eta}_{t}m_{t-1}\frac{(1-\sqrt{\beta_{2}})\epsilon}{\sqrt{\beta_{2}v_{t-1}}+\sqrt{\beta_{2}}\epsilon}+\beta_{1}\hat{\eta}_{t}m_{t-1}\frac{(1-\sqrt{\beta_{2}})(1-\beta_{2})\epsilon(\sigma_{t}^{2}-\hat{g}_{t}^{2})}{(\sqrt{v_{t}}+\epsilon)(\sqrt{v_{t}}+\sqrt{\hat{v}_{t}})(\sqrt{\beta_{2}v_{t-1}}+\sqrt{\beta_{2}}\epsilon)}.\end{split}$
For simplicity, we denote
[…]
Then, we obtain
$\begin{split}\Delta_{t}-\frac{\beta_{1}\alpha_{t}}{\sqrt{\beta_{2}}\alpha_{t-1}}\Delta_{t-1}&=-\frac{\alpha_{t}m_{t}}{\sqrt{v_{t}}+\epsilon}+\frac{\beta_{1}\alpha_{t}m_{t-1}}{\sqrt{\beta_{2}v_{t-1}+\sqrt{\beta_{2}}\epsilon}}\\ &=-\frac{(1-\beta_{1})\alpha_{t}\hat{g}_{t}}{\sqrt{v_{t}}+\epsilon}+\beta_{1}\alpha_{t}m_{t-1}\left(\frac{1}{\sqrt{\beta_{2}v_{t-1}}+\sqrt{\beta_{2}}\epsilon}-\frac{1}{\sqrt{v_{t}}+\epsilon}\right)\\ &=-(1-\beta_{1})\hat{\eta}_{t}\hat{g}_{t}-\sqrt{\hat{\eta}_{t}}A_{t}^{1}+\sqrt{\hat{\eta}_{t}}A_{t}^{2}+\sqrt{\hat{\eta}_{t}}A_{t}^{3}+\sqrt{\hat{\eta}_{t}}A_{t}^{4}+\sqrt{\hat{\eta}_{t}}A_{t}^{5}\end{split}$
Thus, it holds that
$\begin{split}\mathbb{E}\left\langle\nabla f(\theta_{t}),\Delta_{t}\right\rangle&=\frac{\beta_{1}\alpha_{t}}{\sqrt{\beta_{2}}\alpha_{t-1}}\left\langle\nabla f(\theta_{t}),\Delta_{t-1}\right\rangle+\mathbb{E}\left\langle\nabla f(\theta_{t}),\Delta_{t}-\frac{\beta_{1}\alpha_{t}}{\sqrt{\beta_{2}}\alpha_{t-1}}\Delta_{t-1}\right\rangle\\ &=\frac{\beta_{1}\alpha_{t}}{\sqrt{\beta_{2}}\alpha_{t-1}}\left(\mathbb{E}\langle\nabla f(\theta_{t}),\Delta_{t-1}\rangle+\mathbb{E}\langle\nabla f(\theta_{t})-\nabla f(\theta_{t-1}),\Delta_{t-1}\rangle\right)\\ &\qquad+\mathbb{E}\langle\nabla f(\theta_{t}),-(1-\beta_{1})\hat{\eta}_{t}\hat{g}_{t}\rangle+\mathbb{E}\langle\nabla f(\theta_{t}),-\sqrt{\hat{\eta}_{t}}A_{t}^{1}\rangle+\mathbb{E}\langle\nabla f(\theta_{t}),\sqrt{\hat{\eta}_{t}}A_{t}^{2}\rangle\\ &\qquad+\mathbb{E}\langle\nabla f(\theta_{t}),\sqrt{\hat{\eta}_{t}}A_{t}^{3}\rangle+\mathbb{E}\langle\nabla f(\theta_{t}),\sqrt{\hat{\eta}_{t}}A_{t}^{4}\rangle+\mathbb{E}\langle\nabla f(\theta_{t}),\sqrt{\hat{\eta}_{t}}A_{t}^{5}\rangle\end{split}$ (6)
For the first term of (6), it holds that
$\begin{split}&\frac{\beta_{1}\alpha_{t}}{\sqrt{\beta_{2}}\alpha_{t-1}}\left(\mathbb{E}\langle\nabla f(\theta_{t}),\Delta_{t-1}\rangle+\mathbb{E}\langle\nabla f(\theta_{t})-\nabla f(\theta_{t-1}),\Delta_{t-1}\rangle\right)\\ &\leq\frac{\beta_{1}\alpha_{t}}{\sqrt{\beta_{2}}\alpha_{t-1}}\left(\mathbb{E}\langle\nabla f(\theta_{t}),\Delta_{t-1}\rangle+\mathbb{E}\|\nabla f(\theta_{t})-\nabla f(\theta_{t-1})\|\|\Delta_{t-1}\|\right)\\ &\leq\frac{\beta_{1}\alpha_{t}}{\sqrt{\beta_{2}}\alpha_{t-1}}\left(\mathbb{E}\langle\nabla f(\theta_{t}),\Delta_{t-1}\rangle+L\mathbb{E}\|\Delta_{t-1}\|^{2}\right)\\ &=\frac{\beta_{1}\alpha_{t}}{\sqrt{\beta_{2}}\alpha_{t-1}}M_{t-1}\end{split}$
For the second term of (6), it holds that
$\mathbb{E}\langle\nabla f(\theta_{t}),-(1-\beta_{1})\hat{\eta}_{t}\hat{g}_{t}\rangle\leq-(1-\beta_{1})h_{0}\mathbb{E}\|\nabla f(\theta_{t})\|_{\hat{\eta}_{t}}^{2}.$
For the rest of the terms, it holds that
[…]
On the other hand, it holds that
$\begin{split}&\left\|A_{t}^{1}\right\|^{2}\leq(1-\beta_{1})^{2}\frac{4\alpha\beta G^{4}}{T\epsilon^{3}},\left\|A_{t}^{2}\right\|^{2}\leq\beta_{1}^{2}\frac{\alpha\beta G^{4}}{T\beta_{2}\epsilon^{3}},\left\|A_{t}^{3}\right\|^{2}\leq\beta_{1}^{2}\frac{\alpha\beta G^{2}}{(1-\zeta)\epsilon T\beta_{2}},\\ &\left\|A_{t}^{4}\right\|^{2}\leq\beta_{1}^{2}\frac{\alpha\beta G^{4}}{T\beta_{2}},\left\|A_{t}^{5}\right\|^{2}\leq\beta_{1}^{2}\frac{\alpha\beta G^{2}}{(1-\zeta)\beta_{2}T}\end{split}$
Define $N_{t}=\frac{C_{2}}{T}+L\mathbb{E}\|\Delta_{t}\|^{2}$, where $C_{2}=\frac{5}{2(1-\beta_{1})h_{0}}\left((1-\beta_{1})^{2}\frac{4\alpha\beta G^{4}}{\epsilon^{3}}+\beta_{1}^{2}\alpha\beta\left(\frac{G^{4}}{\beta_{2}\epsilon^{3}}+\frac{(1+\epsilon)G^{2}}{(1-\zeta)\epsilon\beta_{2}}+\frac{G^{4}}{\beta_{2}}\right)\right)$. It holds that
$M_{t}\leq\frac{\beta_{1}\alpha_{t}}{\sqrt{\beta_{2}}\alpha_{t-1}}M_{t-1}+N_{t}-\frac{1-\beta_{1}}{2}\hat{\eta}_{t}\mathbb{E}\|\nabla f(\theta_{t})\|^{2}_{\hat{\eta}_{t}}\leq\sum_{i=1}^{t}\sqrt{\zeta}^{t-i}N_{i}-\frac{1-\beta_{1}}{2}h_{0}\mathbb{E}\|\nabla f(\theta_{t})\|_{\hat{\eta}_{t}}^{2}$
Thus, by summing $t$ from 1 to $T$, it holds that
$\begin{split}\sum_{t=1}^{T}M_{t}&\leq\sum_{t=1}^{T}\sum_{i=1}^{t}\sqrt{\zeta}^{t-i}N_{i}-\frac{(1-\beta_{1})h_{0}}{2}\mathbb{E}\|\nabla f(\theta_{t})\|_{\hat{\eta}_{t}}^{2}\\ &\leq\frac{1}{1-\sqrt{\zeta}}\sum_{t=1}^{T}N_{t}-\frac{(1-\beta_{1})h_{0}}{2}\mathbb{E}\|\nabla f(\theta_{t})\|_{\hat{\eta}_{t}}^{2}\\ &\leq\frac{C_{2}}{1-\sqrt{\zeta}}+\frac{LG^{2}\alpha^{2}}{(1-\sqrt{\zeta})\epsilon}-\frac{(1-\beta_{1})h_{0}}{2}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\theta_{t})\|_{\hat{\eta}_{t}}^{2}.\end{split}$
Lemma F.11.
Let $\tau$ be randomly chosen from $\{1,2,\cdots,T\}$ with equal probabilities $p_{\tau}=\frac{1}{T}$. We have the following estimate:
$\mathbb{E}[\|\nabla f\left(\theta_{\tau}\right)\|^{2}]\leq\frac{\sqrt{G^{2}+\epsilon d}}{\alpha\sqrt{T}}\mathbb{E}\left[\sum_{t=1}^{T}\|\nabla f\left(\theta_{t}\right)\|_{\hat{\eta_{t}}}^{2}\right].$
Proof.
Note that $\|\hat{v}_{t}\|_{1}=\beta_{2}\|v_{t-1}\|_{1}+\left(1-\beta_{2}\right)\|\sigma_{t}\|^{2}$ and $\|\hat{g}_{t}\|\leq G$. It is straightforward to prove $\|v_{t}\|_{1}\leq G^{2}$. Hence, we have $\|\hat{v}_{t}+\epsilon\|_{1}\leq G^{2}+\epsilon d$.
Utilizing this inequality, we have
$\begin{split}\|\nabla f\left(\theta_{t}\right)\|^{2}&=\frac{\|\nabla f\left(\theta_{t}\right)\|^{2}}{\sqrt{\|\hat{v}_{t}+\epsilon\|_{1}}}\sqrt{\|\hat{v}_{t}+\epsilon\|_{1}}=\sqrt{\|\hat{v}_{t}+\epsilon\|_{1}}\sum_{k=1}^{d}\frac{|\nabla_{k}f\left(\theta_{t}\right)|^{2}}{\sqrt{\sum_{l=1}^{d}\hat{v}_{t,l}+\epsilon}}\\ &\leq\sqrt{\|\hat{v}_{t}+\epsilon\|_{1}}\alpha_{t}^{-1}\sum_{k=1}^{d}\frac{\alpha_{t}}{\sqrt{\hat{v}_{t,k}+\epsilon}}|\nabla_{k}f\left(\theta_{t}\right)|^{2}=\sqrt{\|\hat{v}_{t}+\epsilon\|_{1}}\alpha_{t}^{-1}\|\nabla f\left(\theta_{t}\right)\|_{\hat{\eta}_{t}}^{2}\\ &\leq\sqrt{G^{2}+\epsilon d}\alpha_{t}^{-1}\|\nabla f\left(\theta_{t}\right)\|_{\hat{\eta}_{t}}^{2}\leq\frac{\sqrt{G^{2}+\epsilon d}}{\alpha_{T}}\|\nabla f\left(\theta_{t}\right)\|^{2}_{\hat{\eta}_{t}}.\end{split}$
Then, by using the definition of $\theta_{\tau}$, we obtain
$\mathbb{E}\left[\|\nabla f\left(\theta_{\tau}\right)\|^{2}\right]=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[\|\nabla f\left(\theta_{t}\right)\|^{2}\right]\leq\frac{\sqrt{G^{2}+\epsilon d}}{\alpha\sqrt{T}}\mathbb{E}\left[\sum_{t=1}^{T}\|\nabla f\left(\theta_{t}\right)\|^{2}_{\hat{\eta}_{t}}\right].$
Thus, the desired result is obtained. ∎
Theorem F.12.
Let $\{\theta_{t}\}$ be a sequence generated by AdaGC for initial values $\theta_{1}$ and $m_{0}=v_{0}=0$. Assumptions F.2 to F.5 hold. With the hyperparameters $\alpha_{t}=\alpha/\sqrt{T}$$\beta_{2}=1-\beta/T$ and $\zeta=\beta_{1}^{2}/\beta_{2}<1$. Let $\tau$ be randomly chosen from $\{1,2,\cdots,T\}$ with equal probabilities. We have
$\mathbb{E}\|\nabla f(\theta_{\tau})\|^{2}\leq\frac{C}{\sqrt{T}}$
where $C=\frac{\sqrt{G^{2}+\epsilon d}}{\alpha}\left(f(\theta_{1})-\underline{f}+\frac{C_{2}}{1-\sqrt{\zeta}}+\frac{LG^{2}\alpha^{2}}{(1-\sqrt{\zeta})\epsilon}\right)$ and $C_{2}=\frac{5}{2(1-\beta_{1})h_{0}}\left((1-\beta_{1})^{2}\frac{4\alpha\beta G^{4}}{\epsilon^{3}}+\beta_{1}^{2}\alpha\beta\left(\frac{G^{4}}{\beta_{2}\epsilon^{3}}+\frac{(1+\epsilon)G^{2}}{(1-\zeta)\epsilon\beta_{2}}+\frac{G^{4}}{\beta_{2}}\right)\right)$.
Proof.
With the Lipschitz continuity condition of $f$, it holds that
$\mathbb{E}f(\theta_{t+1})\leq\mathbb{E}\left[f(\theta_{t})+\langle\nabla f(\theta_{t}),\Delta_{t}\rangle+\frac{L}{2}\|\Delta_{t}\|^{2}\right]\leq\mathbb{E}f(\theta_{t})+M_{t}.$
By summing $t$ from $1$ to $T$, it holds that
$\mathbb{E}f(\theta_{T+1})\leq f(\theta_{1})+\sum_{t=1}^{T}M_{t}\leq f(\theta_{1})+\frac{C_{2}}{1-\sqrt{\zeta}}+\frac{LG^{2}\alpha^{2}}{(1-\sqrt{\zeta})\epsilon}-\frac{(1-\beta_{1})h_{0}}{2}\sum_{t=1}^{T}\mathbb{E}\|\nabla f(\theta_{t})\|_{\hat{\eta}_{t}}^{2}$
Thus, it holds that
$\begin{split}\mathbb{E}\left[\|\nabla f(\theta_{\tau}\|^{2}\right]&\leq\frac{\sqrt{G^{2}+\epsilon d}}{\alpha\sqrt{T}}\mathbb{E}\left[\sum_{t=1}^{T}\|\nabla f(\theta_{t})\|_{\hat{\eta}_{t}}^{2}\right]\\ &\leq\frac{\sqrt{G^{2}+\epsilon d}}{\alpha\sqrt{T}}\left(f(\theta_{1})-\mathbb{E}[f(\theta_{T+1})]+\frac{C_{2}}{1-\sqrt{\zeta}}+\frac{LG^{2}\alpha^{2}}{(1-\sqrt{\zeta})\epsilon}\right)\\ \ &\leq\frac{\sqrt{G^{2}+\epsilon d}}{\alpha\sqrt{T}}\left(f(\theta_{1})-\underline{f}+\frac{C_{2}}{1-\sqrt{\zeta}}+\frac{LG^{2}\alpha^{2}}{(1-\sqrt{\zeta})\epsilon}\right)\end{split}$